a rectangle has perimeter 56cm. if 4 cm is taken from the length and added to the width, the rectangle becomes a square. Find the dimensions of the original rectangle.
Answer by W3L(1)

Let length and width = L and W.
from statement "A rectangle has perimeter 56 cm. If 4 cm is taken from the length and added to the width, the rectangle becomes a square", you can deduce
2L + 2W = 56 ( its sides make up its perimeter )
L - 4 = W + 4
L = W + 8
sub L = W + 8 into 2L + 2W = 56
2 ( W + 8 ) + 2W = 56
2W + 16 + 2W = 56
4W = 40
W = 10
sub W = 10 into L = W + 8
L = 10 + 8 = 18
original dimensions of rectangle, length = 18cm and width = 10cm